OpenStudy (anonymous):

Determine the limit, if it exists, of the sequences whose nth term is given a_n = (-1)^n * (n/(n+1))

OpenStudy (anonymous):

limit as \(n\to \infty\) ?

OpenStudy (anonymous):

yes, i'm sorry if i left that out.

OpenStudy (anonymous):

I was thinking of taking the abs value and finding the limit, but upon doing that i get an answer of 1 which I know I cant use it if it does not equal 1.

OpenStudy (anonymous):

if you did not have that annoying \((-1)^n\) you would take \[\lim_{n\to \infty}\frac{n}{n+1}=1\] but the term \((-1)^n\) makes it oscillate so no limit goes to one and minus one

OpenStudy (anonymous):

*if it does not equal 0

OpenStudy (anonymous):

in other words it has two limits, which means it has no limit, but rather a lim inf and a lim sup

OpenStudy (anonymous):

so when doing a problem like this would I always take out the -1^n and evaluate ? even though the theorem says it would have to equal 0 for that limit to be the same ?

OpenStudy (anonymous):

nvm. I get it I missed the part where you say it oscillates. Thanks!